OpenStudy (awesome781):

How can I factor the following equation: 28n^4+16n^3-80n^2

OpenStudy (amoodarya):

first factor n^2

OpenStudy (amoodarya):

then factor 4

OpenStudy (amoodarya):

4n^2 (7n^2-4n-20)

OpenStudy (awesome781):

so I have to distribute?...the 4n^2 to the numbers in the parenthesis?

OpenStudy (phi):

no, factoring is the opposite of distributing

OpenStudy (awesome781):

ok

OpenStudy (phi):

the next step is to see if you can factor 7n^2-4n-20

OpenStudy (awesome781):

oh...ok

OpenStudy (awesome781):

What if you can't?

OpenStudy (phi):

you don't pass ?

OpenStudy (phi):

that 7 in front of the n^2 makes factoring this a pain. one way to factor it is first multiply 7*20 (the 7 times the last number, ignoring the sign) you get 140 now we need to list pairs of numbers that multiply to give 140 1,140 2,70 4,35 5,28 7,20 10,14 and look for a pair with *opposite signs* add up to -4 (the middle number in your problem)

OpenStudy (awesome781):

the last pair

OpenStudy (phi):

the last pair work, if we use +10 and -14 we know the 7 in front of n^2 means we will have (7n )(n ) we know we need a -14 to get a -14 (using that 7) is put -2 in the 2nd parens

OpenStudy (awesome781):

yes exactly

OpenStudy (phi):

that leaves a +10 for the (7n )

OpenStudy (phi):

(7n+10)(n-2) I would check this answer, by multiplying it out: 7n*n = 7n^2 7n* -2= -14n 10*n= 10n 10*-2 = -20 add up 7n^2 -14n + 10n -20 simplify 7n^2 -4n -20

OpenStudy (awesome781):

Alright.

OpenStudy (awesome781):

Ty :D phi